OpenStudy (anonymous):

A surfboard cost 150 dollars you have a 10% off on the sale price of a surfboard. Which is the better buy. Is 40% off the regular price better or is 30% off the regular price and than 10% off the sale price

OpenStudy (study312):

40% as a decimal is 0.40 so 0.40 • 150 = x then 150 - x = y. now choice B is 30% 0.30, so 0.30 • 150 = x, 150 - x = y, 10% is 0.10 so 0.10 • y = z. y - z = your answer.

OpenStudy (anonymous):

@Study312 Okay thanks I'll try that